xx Z4c Thorn for use with CarpetX

Z4c Thorn for use with CarpetX

Erik Schnetter <eschnetter@perimeterinstitute.ca >

April 26, 2024

Abstract

This Thorn implements the Z4c formulation of the Einstein Equationsx[?]. It also monitors the stability of those equationsx[?].

1 Using This Thorn

The SpacetimeX/Z4c thorn reads (i.e. in Z4cxInitial1) and, if the parameter Z4c::calcxADMxvars is set

(which is true by default), writes (i.e. in Z4cxADM) the ADM variables (i.e. CarpetX/ADMBaseX) automatically. Derivatives are hard-coded to fourth order. However, should it be desired, second order is also possible by editing the constant derivxorder in derivs.hxx to the value 2. This thorn integrates with the CarpetX/ODESolvers thorn.

References

@articleHilditch:2012fp, author = ”Hilditch, David and Bernuzzi, Sebastiano and Thierfelder, Marcus and Cao, Zhoujian and Tichy, Wolfgang and Bruegmann, Bernd”, title = ”Compact binary evolutions with the Z4c formulation”, journal = ”Phys. Rev.”, volume = ”D88”, year = ”2013”, pages = ”084057”, doi = ”10.1103/PhysRevD.88.084057”, eprint = ”1212.2901”, archivePrefix = ”arXiv”, primaryClass = ”gr-qc”, SLACcitation = ” @articlecao2012numerical, title=Numerical stability of the Z4c formulation of general relativity, author=Cao, Zhoujian and Hilditch, David, journal=Physical Review D, volume=85, number=12, pages=124032, year=2012, publisher=APS

2 Parameters

11 alphagxfloor 10 Scope: private01 REAL

11 Description: Floor for alphaG

11  Range  10              01  Default: 1.0e-10

11 (0:*  10

11

11 boundaryxconditions10 Scope: private01  KEYWORD

11 Description: boundary conditions

11 Range 10 01 Default: CarpetX
11 NewRadX  10 radiative boundary conditions using NewRadX
11 CarpetX  10 use CarpetX default boundary conditions

11

11 calcxadmxvars  10 Scope: private01  BOOLEAN

11 Description: Calculate ADM variables

11                   10              01  Default: yes

11

11 calcxadmrhsxvars 10 Scope: private01  BOOLEAN

11 Description: Calculate RHS of ADM variables

11                   10              01  Default: yes

11

11 calcxconstraints  10 Scope: private01  BOOLEAN

11 Description: Calculate constraints

11                   10              01  Default: yes

11 chixfloor  10 Scope: private01  REAL

11 Description: Floor for chi

11  Range  10              01  Default: 1.0e-10

11 (0:*  10

11

11 epsdiss  10 Scope: private01  REAL

11 Description: Dissipation coefficient <arXiv:gr-qc/0610128>

11 Range 10 01 Default: 0.32
11 0.0:*  10

11

11 eta  10 Scope: private01  REAL

11 Description: eta

11  Range  10              01  Default: 2.0

11 *:*  10

11

11 fxmuxl  10 Scope: private01  REAL

11 Description: muxL = fxmuxL / alpha

11  Range  10              01  Default: 2.0

11 *:*  10

11

11 fxmuxs  10 Scope: private01  REAL

11 Description: muxS = fxmuxS / alphaˆ2

11  Range  10              01  Default: 1.0

11 *:*  10

11

11 kappa1  10 Scope: private01  REAL

11 Description: kappa1

11  Range  10              01  Default: 0.02

11 *:*  10

11

11 kappa2  10 Scope: private01  REAL

11 Description: kappa2

11  Range  10              01  Default: 0.0

11 *:*  10

11

11 nxalphag  10 Scope: private01  INT

11 Description: n power of outgoing boundary rn̂ fall off rate for alpha

11 Range 10 01 Default: 1
11 0:2  10 1 is my guess

11

11 nxat  10 Scope: private01  INT

11 Description: n power of outgoing boundary rn̂ fall off rate for Atxij

11 Range 10 01 Default: 1
11 0:2  10 Maybe 1?

11 nxbetag  10 Scope: private01  INT

11 Description: n power of outgoing boundary rn̂ fall off rate for beta

11 Range 10 01 Default: 1
11 0:2  10 1 is my guess

11

11 nxchi  10 Scope: private01  INT

11 Description: n power of outgoing boundary rn̂ fall off rate for chi

11 Range 10 01 Default: 1
11 0:2  10 1 is reasonable

11

11 nxgammat  10 Scope: private01  INT

11 Description: n power of outgoing boundary rn̂ fall off rate for gammatxij

11 Range 10 01 Default: 1
11 0:2  10 1 is reasonable

11

11 nxgamt  10 Scope: private01  INT

11 Description: n power of outgoing boundary rn̂ fall off rate for Gamtx

11 Range 10 01 Default: 1
11 0:2  10 Maybe 1?

11

11 nxkh  10 Scope: private01  INT

11 Description: n power of outgoing boundary rn̂ fall off rate for Kh

11 Range 10 01 Default: 1
11 0:2  10 Maybe 1?

11

11 nxtheta  10 Scope: private01  INT

11 Description: n power of outgoing boundary rn̂ fall off rate for Theta

11 Range 10 01 Default: 1
11 0:2  10 Maybe 1?

11

11 setxthetaxzero  10 Scope: private01  BOOLEAN

11 Description: set Theta to zero, which converts Z4c to BSSN

11                   10              01  Default: no

11

3 Interfaces General Implements: z4c

Inherits: admbasex tmunubasex

Grid Variables

3.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 chi 10 chi 01 compact 11 0
11                    10                    01 description 11 chi
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”chixrhs” dependents=”ADMBaseX::metric”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gammaxtilde  10                    01  compact 11 0
11                    10  gammatxx  01 description 11 gamma-tilde
11                    10  gammatxy  01 dimensions 11 3
11                    10  gammatxz  01 distribution 11 DEFAULT
11                    10  gammatyy  01 group type 11 GF
11                    10  gammatyz  01  tags 11 parities={+1 +1 +1 -1 -1 +1 -1 +1 -1 +1 +1 +1 +1 -1 -1 +1 +1 +1} rhs=”gammaxtildexrhs” dependents=”ADMBaseX::metric”
11                    10  gammatzz  01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  kxhat  10                    01  compact 11 0
11                    10  Kh  01 description 11 K-hat
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”Kxhatxrhs” dependents=”ADMBaseX::curv”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  axtilde  10                    01  compact 11 0
11                    10  Atxx  01 description 11 A-tilde
11                    10  Atxy  01 dimensions 11 3
11                    10  Atxz  01 distribution 11 DEFAULT
11                    10  Atyy  01 group type 11 GF
11                    10  Atyz  01  tags 11 parities={+1 +1 +1 -1 -1 +1 -1 +1 -1 +1 +1 +1 +1 -1 -1 +1 +1 +1} rhs=”Axtildexrhs” dependents=”ADMBaseX::curv”
11                    10  Atzz  01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gamxtilde  10                    01  compact 11 0
11                    10  Gamtx  01 description 11 Gamma-tilde
11                    10  Gamty  01 dimensions 11 3
11                    10  Gamtz  01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 parities={-1 +1 +1 +1 -1 +1 +1 +1 -1} rhs=”Gamxtildexrhs”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 theta 10 Theta 01 compact 11 0
11                    10                    01 description 11 Theta
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”Thetaxrhs” dependents=”ADMBaseX::curv”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 alphag 10 alphaG 01 compact 11 0
11                    10                    01 description 11 alpha
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”alphaGxrhs” dependents=”ADMBaseX::lapse ADMBaseX::dtlapse”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  betag  10                    01  compact 11 0
11                    10  betaGx  01 description 11 beta
11                    10  betaGy  01 dimensions 11 3
11                    10  betaGz  01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 parities={-1 +1 +1 +1 -1 +1 +1 +1 -1} rhs=”betaGxrhs” dependents=”ADMBaseX::shift ADMBaseX::dtshift”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  ztc  10                    01  compact 11 0
11                    10  ZtCx  01 description 11 Z-tilde
11                    10  ZtCy  01 dimensions 11 3
11                    10  ZtCz  01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 parities={-1 +1 +1 +1 -1 +1 +1 +1 -1} checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 hc 10 HC 01 compact 11 0
11                    10                    01 description 11 H
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  mtc  10                    01  compact 11 0
11                    10  MtCx  01 description 11 M-tilde
11                    10  MtCy  01 dimensions 11 3
11                    10  MtCz  01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 parities={-1 +1 +1 +1 -1 +1 +1 +1 -1} checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 allc 10 allC 01 compact 11 0
11                    10                    01 description 11 constraint monitor
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 chixrhs 10 chixrhs 01 compact 11 0
11                    10                    01 description 11 chi
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 gammaxtildexrhs 10                    01  compact 11 0
11                    10  gammatxxxrhs 01 description 11 gamma-tilde
11                    10  gammatxyxrhs 01 dimensions 11 3
11                    10  gammatxzxrhs 01 distribution 11 DEFAULT
11                    10  gammatyyxrhs 01 group type 11 GF
11 10 gammatyzxrhs 01 tags 11 parities={+1 +1 +1 -1 -1 +1 -1 +1 -1 +1 +1 +1 +1 -1 -1 +1 +1 +1} checkpoint=”no”
11                    10  gammatzzxrhs 01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  kxhatxrhs  10                    01  compact 11 0
11                    10  Khxrhs  01 description 11 K-hat
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  axtildexrhs  10                    01  compact 11 0
11                    10  Atxxxrhs  01 description 11 A-tilde
11                    10  Atxyxrhs  01 dimensions 11 3
11                    10  Atxzxrhs  01 distribution 11 DEFAULT
11                    10  Atyyxrhs  01 group type 11 GF
11                    10  Atyzxrhs  01  tags 11 parities={+1 +1 +1 -1 -1 +1 -1 +1 -1 +1 +1 +1 +1 -1 -1 +1 +1 +1} checkpoint=”no”
11                    10  Atzzxrhs  01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gamxtildexrhs  10                    01  compact 11 0
11                    10  Gamtxxrhs  01 description 11 Gamma-tilde
11                    10  Gamtyxrhs  01 dimensions 11 3
11                    10  Gamtzxrhs  01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 parities={-1 +1 +1 +1 -1 +1 +1 +1 -1} checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 thetaxrhs 10 Thetaxrhs 01 compact 11 0
11                    10                    01 description 11 Theta
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 alphagxrhs 10 alphaGxrhs 01 compact 11 0
11                    10                    01 description 11 alpha
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  betagxrhs  10                    01  compact 11 0
11                    10  betaGxxrhs  01 description 11 beta
11                    10  betaGyxrhs  01 dimensions 11 3
11                    10  betaGzxrhs  01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 parities={-1 +1 +1 +1 -1 +1 +1 +1 -1} checkpoint=”no”
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

Uses header: defs.hxx div.hxx dual.hxx loopxdevice.hxx mat.hxx simd.hxx sum.hxx vec.hxx vect.hxx newradx.hxx

4 Schedule This section lists all the variables which are assigned storage by thorn SpacetimeX/Z4c. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 Always:  11 x

11 chi  11 x

11 gammaxtilde  11 x

11 MtC  11 x

11 allC  11 x
11 chixrhs  11 x
11 gammaxtildexrhs 11 x
11 Kxhatxrhs  11 x
11 Axtildexrhs  11 x
11 Gamxtildexrhs 11 x
11 Thetaxrhs  11 x
11 alphaGxrhs  11 x
11 betaGxrhs  11 x
11 Kxhat  11 x
11 Axtilde  11 x
11 Gamxtilde  11 x
11 Theta  11 x
11 alphaG  11 x
11 betaG  11 x

11 ZtC  11 x
11 HC  11 x
11 x  11 x

11

Scheduled Functions CCTKxINITIAL x z4cxinitialgroup x convert adm to z4c variables x

11 x 11 After: 11 admbasexxpostinitial
11 x 11 Type: 11 group

11

CCTKxINITIAL (conditional) x z4cxpoststepgroup x post-process z4c variables x

11 x 11 After: 11 z4cxinitialgroup
11 x 11 Before: 11 admbasexxsetadmvars
11 x 11 Type: 11 group

11

Z4cxInitialGroup x z4cxinitial1 x convert adm to z4c variables, part 1 x

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbasex::metric(interior)
11 x 11 x  11 admbasex::curv(interior)
11 x 11 x  11 admbasex::lapse(interior)
11 x 11 x  11 admbasex::shift(interior)
11 x 11 Sync:  11 gammaxtilde
11 x 11 Type:  11 function
11 x 11 Writes: 11 chi(interior)
11 x 11 x  11 gammaxtilde(interior)
11 x 11 x  11 kxhat(interior)
11 x 11 x  11 axtilde(interior)
11 x 11 x  11 theta(interior)
11 x 11 x  11 alphag(interior)
11 x 11 x  11 betag(interior)

11

Z4cxInitialGroup x z4cxinitial2 x convert adm to z4c variables, part 2 x

11 x 11 After:  11 z4cxinitial1
11 x 11 Language: 11 c
11 x 11 Reads:  11 gammaxtilde(everywhere)
11 x 11 Type:  11 function
11 x 11 Writes: 11 gamxtilde(interior)

11

Z4cxPostStepGroup x z4cxenforce x enforce algebraic z4c constraints

11 x 11 Language: 11 c
11 x 11 Reads:  11 chi(interior)
11 x 11 x  11 gammaxtilde(interior)
11 x 11 x  11 axtilde(interior)
11 x 11 x  11 alphag(interior)
11 x 11 Sync:  11 chi
11 x 11 x  11 gammaxtilde
11 x 11 x  11 kxhat
11 x 11 x  11 axtilde
11 x 11 x  11 gamxtilde
11 x 11 x  11 theta
11 x 11 x  11 alphag
11 x 11 x  11 betag
11 x 11 Type:  11 function
11 x 11 Writes: 11 chi(interior)
11 x 11 x  11 gammaxtilde(interior)
11 x 11 x  11 axtilde(interior)
11 x 11 x  11 alphag(interior)

11

Z4cxPostStepGroup (conditional) x z4cxadm x convert z4c to adm variables x

11 x 11 After:  11 z4cxenforce
11 x 11 Language: 11 c
11 x 11 Reads:  11 chi(everywhere)
11 x 11 x  11 gammaxtilde(everywhere)
11 x 11 x  11 kxhat(everywhere)
11 x 11 x  11 axtilde(everywhere)
11 x 11 x  11 gamxtilde(everywhere)
11 x 11 x  11 theta(everywhere)
11 x 11 x  11 alphag(everywhere)
11 x 11 x  11 betag(everywhere)
11 x 11 Type:  11 function
11 x 11 Writes: 11 admbasex::metric(everywhere)
11 x 11 x  11 admbasex::curv(everywhere)
11 x 11 x  11 admbasex::lapse(everywhere)
11 x 11 x  11 admbasex::dtlapse(everywhere)
11 x 11 x  11 admbasex::shift(everywhere)
11 x 11 x  11 admbasex::dtshift(everywhere)

11

Z4cxPostStepGroup2 (conditional) x z4cxadm2 x calculate second time derivatives of adm variables

11 x 11 Language: 11 c
11 x 11 Reads:  11 chi(everywhere)
11 x 11 x  11 gammaxtilde(everywhere)
11 x 11 x  11 kxhat(everywhere)
11 x 11 x  11 axtilde(everywhere)
11 x 11 x  11 gamxtilde(everywhere)
11 x 11 x  11 theta(everywhere)
11 x 11 x  11 alphag(everywhere)
11 x 11 x  11 betag(everywhere)
11 x 11 x  11 tmunubasex::ettt(interior)
11 x 11 x  11 tmunubasex::etti(interior)
11 x 11 x  11 tmunubasex::etij(interior)
11 x 11 Sync:  11 admbasex::dtcurv
11 x 11 x  11 admbasex::dt2lapse
11 x 11 x  11 admbasex::dt2shift
11 x 11 Type:  11 function
11 x 11 Writes: 11 admbasex::dtcurv(interior)
11 x 11 x  11 admbasex::dt2lapse(interior)
11 x 11 x  11 admbasex::dt2shift(interior)

11

Z4cxAnalysisGroup (conditional) x z4cxconstraints x calculate z4c constraints x

11 x 11 Language: 11 c
11 x 11 Reads:  11 chi(everywhere)
11 x 11 x  11 gammaxtilde(everywhere)
11 x 11 x  11 kxhat(everywhere)
11 x 11 x  11 axtilde(everywhere)
11 x 11 x  11 gamxtilde(everywhere)
11 x 11 x  11 theta(everywhere)
11 x 11 x  11 alphag(everywhere)
11 x 11 x  11 betag(everywhere)
11 x 11 x  11 tmunubasex::ettt(interior)
11 x 11 x  11 tmunubasex::etti(interior)
11 x 11 x  11 tmunubasex::etij(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 ztc(interior)
11 x 11 x  11 hc(interior)
11 x 11 x  11 mtc(interior)
11 x 11 x  11 allc(interior)

11

Z4cxRHSGroup (conditional) x z4cxrhs x calculate z4c rhs

11 x 11 Language: 11 c
11 x 11 Reads:  11 chi(everywhere)
11 x 11 x  11 gammaxtilde(everywhere)
11 x 11 x  11 kxhat(everywhere)
11 x 11 x  11 axtilde(everywhere)
11 x 11 x  11 gamxtilde(everywhere)
11 x 11 x  11 theta(everywhere)
11 x 11 x  11 alphag(everywhere)
11 x 11 x  11 betag(everywhere)
11 x 11 x  11 tmunubasex::ettt(interior)
11 x 11 x  11 tmunubasex::etti(interior)
11 x 11 x  11 tmunubasex::etij(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 chixrhs(interior)
11 x 11 x  11 gammaxtildexrhs(interior)
11 x 11 x  11 kxhatxrhs(interior)
11 x 11 x  11 axtildexrhs(interior)
11 x 11 x  11 gamxtildexrhs(interior)
11 x 11 x  11 thetaxrhs(interior)
11 x 11 x  11 alphagxrhs(interior)
11 x 11 x  11 betagxrhs(interior)

11

Z4cxRHSGroup (conditional) x z4cxapplyxnewradxxboundaryxconditions x apply radiative boundary conditions to z4c rhs variables using newradx x

11 x 11 After:  11 z4cxrhs
11 x 11 Language: 11 c
11 x 11 Reads:  11 chi(everywhere)
11 x 11 x  11 gammaxtilde(everywhere)
11 x 11 x  11 kxhat(everywhere)
11 x 11 x  11 axtilde(everywhere)
11 x 11 x  11 gamxtilde(everywhere)
11 x 11 x  11 theta(everywhere)
11 x 11 x  11 alphag(everywhere)
11 x 11 x  11 betag(everywhere)
11 x 11 x  11 chixrhs(interior)
11 x 11 x  11 gammaxtildexrhs(interior)
11 x 11 x  11 kxhatxrhs(interior)
11 x 11 x  11 axtildexrhs(interior)
11 x 11 x  11 gamxtildexrhs(interior)
11 x 11 x  11 thetaxrhs(interior)
11 x 11 x  11 alphagxrhs(interior)
11 x 11 x  11 betagxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 chixrhs(interior)
11 x 11 x  11 gammaxtildexrhs(interior)
11 x 11 x  11 kxhatxrhs(interior)
11 x 11 x  11 axtildexrhs(interior)
11 x 11 x  11 gamxtildexrhs(interior)
11 x 11 x  11 thetaxrhs(interior)
11 x 11 x  11 alphagxrhs(interior)
11 x 11 x  11 betagxrhs(interior)

11

CCTKxINITIAL x z4cxpoststepgroup2 x post-process z4c variables, part 2

11 x 11 After: 11 tmunubasexxsettmunuvars
11 x 11 x  11 z4cxpoststepgroup
11 x 11 Before: 11 admbasexxsetadmrhs
11 x 11 Type: 11 group

11

CCTKxPOSTREGRID x z4cxpoststepgroup x post-process z4c variables x

11 x 11 Before: 11 admbasexxsetadmvars
11 x 11 Type: 11 group

11

CCTKxPOSTREGRID x z4cxpoststepgroup2 x post-process z4c variables, part 2 x

11 x 11 After: 11 tmunubasexxsettmunuvars
11 x 11 x  11 z4cxpoststepgroup
11 x 11 Before: 11 admbasexxsetadmrhs
11 x 11 Type: 11 group

11

CCTKxANALYSIS x z4cxanalysisgroup x analyse z4c variables x

11 x 11 Type: 11 group

11

ODESolversxPostStep x z4cxpoststepgroup x post-process z4c variables x

11 x 11 Before: 11 admbasexxsetadmvars
11 x 11 Type: 11 group

11

ODESolversxPostStep x z4cxpoststepgroup2 x post-process z4c variables, part 2 x

11 x 11 After: 11 tmunubasexxsettmunuvars
11 x 11 x  11 z4cxpoststepgroup
11 x 11 Before: 11 admbasexxsetadmrhs
11 x 11 Type: 11 group

11

ODESolversxRHS x z4cxrhsgroup x calculate z4c rhs x

11 x 11 Type: 11 group

11

CCTKxWRAGH x z4cxtest x self-test x

11 x 11 Language: 11 c
11 x 11 Options: 11 meta
11 x 11 Type:  11 function